OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 4. Find x when y = 8.

OpenStudy (anonymous):

y=k/3x =>4=k/3(6) =>k=72 =>y=72/3x =>8=72/3x =>x=24/8=3

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

3x = k/y for x = 6, y =4 k = 3xy = 3*6*4 for y = 8 3x = 3*6*4/8 = 9 x = 9/3 = 3 x = 3